OpenStudy (anonymous):

(5,7);y=1/3x+2 ??? (-9,3);3x+y=5 ??? (3,-6);x+y =-4 ??? Need help >_< .

OpenStudy (anonymous):

do the equations be perpendicular with the coordinates?

OpenStudy (anonymous):

"write an equation of the line that passes through the given point and is perpendicular to the graph of the given equation."

OpenStudy (anonymous):

uh, do i have to show all the work or can i just put the answers?

OpenStudy (anonymous):

1. y= -3x + 22 2. y= 1/3x +6 3. y= x-9

OpenStudy (anonymous):

thanks, i really do hate parralel & perpendicular.

OpenStudy (anonymous):

haha it's pretty easy. the parallel lines have the same slopes and the perpendicular lines have the opposite signed reciprocal slopes.